1.     A parallel plate capacitor consists of two rectangular plates, each with an area of  and are separated from each other by a 2.00 mm thick dielectric with a dielectric constant of 5.26. The capacitor is connected to a 12.0 V battery. How much energy is stored in the capacitor?



1
Expert's answer
2021-02-18T18:53:38-0500

Asuume that the area is "0.005\ m^2"


"E=\frac{CU^2}{2}=\frac{\epsilon_0\epsilon AU^2}{2d}=\frac{8.85\cdot10^{-12}\cdot5.26\cdot 0.005\cdot 12^2}{2\cdot 0.002}=8\cdot10^{-9}\ J" . Answer